# Atomic Fact Decomposition Helps Attributed Question Answering

### v1

@@ 4.1.2. Fact decomposition
After obtaining the instruction-tuned Molecule-to-Atomic LLM, ARE decomposes the long-form answer “$X$: The player with the most Super Bowl rings is Tom Brady. Tom Brady is an American football quarterback who has won six Super Bowl championships.” into molecular clauses “$MF_{1}$: The player with the most Super Bowl rings is Tom Brady.” and “$MF_{2}$: Tom Brady is an American football quarterback who has won six Super Bowl championships.”, $MF_{1}$ only has one atomic fact $AF_{11}$, $MF_{2}$ can be further decomposed into two atomic facts “$AF_{21}$: Tom Brady is an American football quarterback.” and “$AF_{22}$: Tom Brady has won six Super Bowl championships.”.
4.2. Atomic Fact based Evidence Retrieval
@@
(2) $\mathbf{e}_{c}^{z},\text{{AF}}_{ij}=\textrm{Sentence-Bert}(e_{c}^{z},AF_{ij}),$
@@
where $e_{c}^{z},z\in\{1,...,o\}$ is the $z$-th searched evidence in $e_{c}$. ARE calculates the relevance score $R$($\mathbf{e}_{c}^{z}$, $\text{{AF}}_{ij}$) through the cosine similarity function:
(3) $R(\mathbf{e}_{c}^{z},\text{{AF}}_{ij})=\frac{\mathbf{e}_{c}^{z}\cdot\text{{AF}}_{ij}}{\parallel\mathbf{e}_{c}^{z}\parallel\parallel\text{{AF}}_{ij}\parallel},$
[…]
4.3. LLM-based Evidence Verifing and Editing
@@
Due to the fact that hallucinated knowledge probably exists in the long-form answers generated by LLMs, ARE utilizes the evidence verifing LLM to assess whether atomic facts need to be edited or re-retrieved through comparing with the retrieved evidence. The evidence verifier takes as input the atomic fact and its corresponding evidence, and outputs different statuses, including:
@@
(4) $\textrm{EV}(AF_{ij},e_{ij})=\left\{\begin{matrix}1,&\textit{supportive}\\ 2,&\textit{editing required}\\ 3,&\textit{irrelevant},\end{matrix}\right.$
@@
where $\text{EV}(\cdot,\cdot)$ denotes the evidence verifier, which determines the relationships between the atomic fact and evidence. Specifically, $1$ represents supportive, requiring no further action; $2$ means editing required, necessitating the revision of atomic facts; $3$ indicates irrelevant, which requires a new evidence retrieval. The prompt template of the evidence verifier can be found in Appendix A.3.
@@
Figure 2 illustrates the process. The status between $AF_{11}$ and $e_{11}$ is classified as supportive, $AF_{21}$ retrieves irrelevant evidence $e_{21}$, thus requiring re-retrieval. $AF_{22}$ needs to be edited based on evidence $e_{22}$. For irrelevant, ARE utilizes the fact expanding LLM to expand the atomic fact $AF_{21}$. The expanded fact contains more complete information, enabling more effective retrieval of relevant evidence. For example, ”Tom Brady” is expanded to ”Thomas Edward Patrick Brady Jr.”. The newly retrieved evidence will be re-verified by the evidence verifier, and this process will be repeated. Details of the prompt template and the effectiveness of fact expansion from the perspective of the information theory can be found in Appendix A.3 and Appendix B.
@@
For editing required, ARE designs prompts for LLMs based on in-context learning and chain-of-thought prompting techniques  (zheng-etal-2023-edit, ; gao2023rarr, ), guiding the LLMs to revise the atomic fact $AF_{22}$ using the retrieved evidence $e_{22}$. Since the edits are made at the atomic fact level rather than revising the entire answer $X$, ARE can precisely adjust the necessary details, thus minimizing unnecessary modification. The detailed editing instructions are shown in Appendix A.4.
@@
The edited atomic facts are re-verified and backtracked to their original positions within the molecular clauses, forming the final revised long-form answer $X^{\prime}$ along with the other molecular clauses that do not require editing. To further obtain the attribution report $A$, all evidences $e_{ij}$ are aggregated into a sequence $E_{i}$ to support the $MF_{i}$. Since each $AF_{ij}$ is derived from the decomposition of $MF_{i}$, there may be overlaps among the evidence $e_{ij}$. To address this issue, duplicate snippets are removed. Ultimately, $A$ is compiled as $\{E_{1},...,E_{i},...,E_{m}\}$.
[…]
Methods GPT-3.5 Llama3-70B Llama2-70B
[…]
@@
DRQA 0.424 0.647 - - - 0.382 0.620 - - - 0.483 0.522 - - - - -
EFEC 0.598 0.042 0.762 0.080 0.670 0.490 0.058 0.717 0.107 0.582 0.357 0.058 0.719 0.107 0.477 0.098 0.576
CCVER 0.624 0.066 0.928 0.123 0.747 0.597 0.071 0.921 0.132 0.724 0.423 0.068 0.813 0.126 0.539 0.127 0.670
RARR 0.649 0.058 0.868 0.109 0.743 0.646 0.060 0.850 0.112 0.734 0.516 0.066 0.594 0.119 0.552 0.113 0.676
ARE 0.670 0.756 0.910 0.826 0.772 0.682 0.739 0.898 0.811 0.759 0.584 0.682 0.926 0.785 0.716 0.807 0.749
@@
DRQA 0.431 0.673 - - - 0.380 0.640 - - - 0.368 0.600 - - - - -
EFEC 0.557 0.040 0.729 0.076 0.632 0.538 0.057 0.728 0.106 0.619 0.498 0.029 0.739 0.056 0.595 0.079 0.615
CCVER 0.630 0.069 0.937 0.129 0.753 0.582 0.073 0.901 0.135 0.707 0.397 0.036 0.850 0.069 0.541 0.111 0.667
RARR 0.646 0.060 0.850 0.112 0.734 0.651 0.065 0.829 0.121 0.729 0.543 0.058 0.679 0.107 0.603 0.113 0.689
ARE 0.716 0.807 0.914 0.857 0.803 0.712 0.767 0.887 0.823 0.790 0.631 0.706 0.940 0.806 0.755 0.829 0.783
@@
DRQA 0.237 0.490 - - - 0.237 0.379 - - - 0.234 0.467 - - - - -
EFEC 0.354 0.049 0.716 0.092 0.474 0.319 0.051 0.666 0.095 0.432 0.361 0.031 0.721 0.059 0.481 0.082 0.462
CCVER 0.372 0.047 0.932 0.089 0.532 0.435 0.063 0.917 0.118 0.590 0.323 0.058 0.854 0.109 0.468 0.105 0.530
RARR 0.356 0.097 0.862 0.174 0.504 0.449 0.073 0.846 0.134 0.586 0.412 0.057 0.604 0.104 0.490 0.138 0.527
ARE 0.463 0.559 0.899 0.689 0.611 0.474 0.502 0.907 0.646 0.623 0.484 0.533 0.912 0.673 0.633 0.667 0.622
@@
DRQA 0.127 0.283 - - - 0.131 0.326 - - - 0.147 0.319 - - - - -
EFEC 0.343 0.071 0.686 0.129 0.457 0.356 0.076 0.698 0.137 0.472 0.357 0.058 0.719 0.107 0.477 0.124 0.469
CCVER 0.292 0.071 0.967 0.132 0.449 0.282 0.081 0.942 0.149 0.434 0.212 0.064 0.845 0.119 0.339 0.133 0.407
RARR 0.340 0.078 0.904 0.144 0.494 0.400 0.084 0.851 0.153 0.544 0.353 0.074 0.606 0.132 0.446 0.143 0.495
ARE 0.412 0.438 0.917 0.593 0.569 0.425 0.417 0.924 0.575 0.582 0.390 0.386 0.942 0.548 0.552 0.572 0.568
5. Experiments
In this section, we outline the experimental setups, present the experimental results, and provide a thorough experimental analysis.
5.1. Evaluation Setups
This section mainly introduces the datasets used and baseline methods for comparison, and the evaluation metrics employed.
@@ 5.1. Evaluation Setups
Dataset. We perform extensive experiments on three Question Answering (QA) datasets: Natural Questions (NQ) (kwiatkowskinatural, ), Mintaka (sen2022mintaka, ), and StrategyQA (geva2021strategyqa, ), as well as the AQA dataset ExpertQA. These datasets are used to evaluate the attribution ability in knowledge-intensive QA task. Following the standard dataset settings (gao2023rarr, ), we randomly select 150 samples from NQ, Mintaka and StrategyQA as test datasets. For ExpertQA, we use the entire provided test set. Details of the datasets can be found in Appendix C.1.
Baselines. The proposed ARE is compared with four post-hoc retrieval based baselines, including: EFEC [^4] https://github.com/j6mes/acl2021-factual-error-correction (thorne-vlachos-2021-evidence, ), DRQA (bohnet2022attributed, ), RARR [^5] https://github.com/anthonywchen/RARR (gao2023rarr, ) and CCVER [^6] https://github.com/jifan-chen/Fact-checking-via-Raw-Evidence (chen-etal-2024-complex, ). More details about the baselines can be found in Appendix C.2.
@@
Metrics. We assess the attribution and editing through several metrics: $Attr_{r}$ is a molecular-level attribution metric, which measures the recall of retrieved evidences (gao2023rarr, ):
(5) $Attr_{r}(X,A)=\underset{MF_{i}\in X}{\textrm{avg}}\ \underset{E_{i}\in A}{\textrm{max}}\ \mathrm{NLI}(E_{i},MF_{i}),$
@@
where, $MF_{i}\in X$ is a molecular clause and $E_{i}\in A$ is the evidence for $MF_{i}$, “max” selects the highest entailment score among all evidences, and “avg” calculates the average score of all evidence, $\mathrm{NLI}(E_{i},MF_{i})$ [^7] https://huggingface.co/google/t5_xxl_true_nli_mixture represents the model probability of $E_{i}$ entailing $MF_{i}$.
@@
However, $Attr_{r}$ focuses solely on evidence recall, neglecting the precision of the recalled evidence. This can result in high scores when the quantity of evidence is sufficiently large. In light of this, we propose $Attr_{p}$ to simultaneously assess the precision and completeness of the evidences by calculating the proportion of invalid evidence among all evidences, which is defined as:
(6) $Attr_{p}=\frac{\sum_{i=1}^{m}\mathbb{I}({\textrm{NLI}_{bi}(E_{i},MF_{i})})}{m},$
where $\mathbb{I}(\mathrm{condition})$ is 1 if the condition is true, 0 for other wise; $E_{i}\in A$ and $MF_{i}\in X$. If $E_{i}$ can not entail $MF_{i}$, it was defined as invalid evidence. As for $\mathrm{NLI}_{bi}(\cdot,\cdot)$ is a binary classification result based on TRUE model. Different from the $\mathrm{NLI}(\cdot,\cdot)$, $\mathrm{NLI}_{bi}(\cdot,\cdot)$ will return true only when the evidence supports the whole sentence, which makes $Attr_{p}$ more stricter than $Attr_{r}$. The proof of $Attr_{p}$ can be used to evaluate the precision is available in Appendix E.
@@
Preservation (gao2023rarr, ) generally utilizes Levenshtein distance to measure the changed information from $X$ to $X^{\prime}$:
(7) $Pres_{(X,X^{\prime})}=\textrm{max}(1-\frac{\textrm{Lev}(X,X^{\prime})}{\textrm{length}(X)},0),$
@@
where $Pres_{(X,X^{\prime})}$ equals 1 when $X$ is identical to $X^{\prime}$, indicating no changes. A value of 0 means $X$ and $X^{\prime}$ share no common words, reflecting complete divergence.
@@
To better compare with different baselines, $F1_{RP}$ (gao2023rarr, ; chen2023purr, ) and $F1_{PP}$ is proposed. $F1_{RP}$ and $F1_{PP}$ are calculated by the following equations:
[…]
More details of the metrics can be found in Appendix C.3.

### v2

@@ IV-A2 Fact decomposition
After obtaining the instruction-tuned Molecule-to-Atomic LLM, ARE decomposes the long-form answer “$X$: The player with the most Super Bowl rings is Tom Brady. Tom Brady is an American football quarterback who has won six Super Bowl championships.” into molecular clauses $MF_{1}$: The player with the most Super Bowl rings is Tom Brady.” and $MF_{2}$: Tom Brady is an American football quarterback who has won six Super Bowl championships.”, $MF_{1}$ only has one atomic fact $AF_{11}$, $MF_{2}$ can be further decomposed into two atomic facts $AF_{21}$: Tom Brady is an American football quarterback.” and $AF_{22}$: Tom Brady has won six Super Bowl championships.”.
IV-B Atomic Fact based Evidence Retrieval
@@
$\mathbf{e}_{c}^{z},\text{{AF}}_{ij}=\textrm{Sentence-Bert}(e_{c}^{z},AF_{ij}),$ (7)
@@
where $e_{c}^{z},z\in\{1,...,o\}$ is the $z$-th searched evidence in $e_{c}$. ARE calculates the relevance score RR($\mathbf{e}_{c}^{z}$, $\text{{AF}}_{ij}$) through the cosine similarity function:
$R(\mathbf{e}_{c}^{z},\text{{AF}}_{ij})=\frac{\mathbf{e}_{c}^{z}\cdot\text{{AF}}_{ij}}{\parallel\mathbf{e}_{c}^{z}\parallel\parallel\text{{AF}}_{ij}\parallel},$ (8)
where $\cdot$” represents the inner product and $\parallel\cdot\parallel$ is the norm of the corresponding vector [3]. Finally, ARE ranks all the evidences $\{e_{c}^{1},...,e_{c}^{o}\}$ of $AF_{ij}$ by relevance score RR($\mathbf{e}_{c}^{z}$, $\text{{AF}}_{ij}$) and selects the top as the most relevant evidence $e_{ij}$ for $AF_{ij}$. The search engine’s response may not always yield valid content, often due to poorly constructed queries or limitations within the search engine itself.
IV-C LLM-based Evidence Verifying and Editing
@@
Given that hallucinated content may exist in long-form answers generated by LLMs, ARE employs an evidence-verifying LLM to assess the factual consistency of atomic facts. This verifier operates using a few-shot, prompt-based approach to determine whether an atomic fact should be retained, edited, or re-retrieved, based on a comparison with the corresponding retrieved evidence. The verifier takes as input an atomic fact along with its associated evidence and outputs one of several status labels, including:
@@
$\textrm{EV}(AF_{ij},e_{ij})=\left\{\begin{matrix}1,&\textit{supportive}\\ 2,&\textit{editing required}\\ 3,&\textit{irrelevant},\end{matrix}\right$. (9)
@@
where $\text{EV}(\cdot,\cdot)$ denotes the evidence verifying LLM, which determines the relationships between the atomic fact and evidence. Specifically, $1$ represents supportive, requiring no further action; $2$ means editing required, necessitating the revision of atomic facts; $3$ indicates irrelevant, which requires a new evidence retrieval. To address potential misjudgments arising from LLM uncertainty, we implement an iterative verification mechanism. This process repeatedly validates results flagged with the statuses editing required and irrelevant, ensuring greater accuracy in the edited or re-retrieved content. The prompt template used by the evidence verifier is provided in Table XI.
Algorithm 1: Atomic fact decomposition-based Retrieval and
Editing framework.
1:
2:
3:
4:
5:
6:
7:
8:
9:
10:
11:
12:
13:
14:
15:
16:
17:
18:
19:
20:
Input: The question q, the corpus of text passages D;
$A\leftarrow$ {};
$X$ $\leftarrow$ LLMs(q);
MFMF, AFAF $\leftarrow$ Fact Decomposition($X$);
for $MF_{i}\in MF$ do
    for $AF_{ij}\in AF$ do
      $e_{ij}$ $\leftarrow$ Evidence Retrieval($AF_{ij}$, $D$);
      for iteration=0: 4 do
          status $\leftarrow$ Evidence Verifying($AF_{ij}$, $e_{ij}$);
          if status == 1 do
            $AF[j]\leftarrow AF_{ij}$
            break
          if status == 2 do
            $AF_{ij}$ $\leftarrow$ Editing($AF_{ij}$, $e_{ij}$)
          if status == 3 do
             Expanded $AF_{ij}$ $\leftarrow$ Fact Expansion($AF_{ij}$)
             $e_{ij}\leftarrow$ Evidence Retrieval(Expanded $AF_{ij}$, $D$);
    AA\ $\cup$ Aggregate($e_{ij}$)
$X^{\prime}\leftarrow$ Backtrack(AFAF, MFMF)
output: The revised answer $X^{\prime}$ and the attribution report AA.
@@
Figure 2 illustrates the process. The status between $AF_{11}$ and $e_{11}$ is classified as supportive, $AF_{21}$ retrieves irrelevant evidence $e_{21}$, thus requiring re-retrieval. $AF_{22}$ needs to be edited based on evidence $e_{22}$. For irrelevant, since LLMs can provide relevant information to guide retrieval systems [48], ARE utilizes the fact expansion LLM to expand the atomic fact $AF_{21}$. The expanded fact contains more complete contextual information, enabling a more effective retrieval of relevant evidence. For example, “Tom Brady” is expanded to “Thomas Edward Patrick Brady Jr.” which is a more complete entity. The newly retrieved evidence $e_{ij}^{\prime}$ will be re-verified by the verifier, and this process will be repeated. The details of the prompt template can be found in Table XI.
@@
For editing required, ARE designs prompts for LLMs based on in-context learning and chain-of-thought prompting techniques  [39, 9], guiding LLMs to revise the atomic fact $AF_{22}$ using the retrieved evidence $e_{22}$. Since the edits are made at the atomic fact level rather than revising the entire answer $X$, ARE can precisely adjust the necessary details, thus minimizing unnecessary modification. Detailed editing instructions are shown in Table XI.
The edited atomic facts are re-verified and reintegrated into their original positions within the molecular clauses, producing the final revised long-form answer $X^{\prime}$. This process also preserves molecular clauses that did not require editing. During backtracking, both molecular and atomic facts are provided in Json format alongside the original long-form answer as a reference. This approach follows the instruction-tuning dataset construction methodology, enabling the LLM to effectively manage interdependencies among atomic facts. The detailed backtracking prompt is presented in Table XI.
@@
To further obtain the attribution report $A$, all evidences $e_{ij}$ are aggregated into a sequence $E_{i}$ to support $MF_{i}$. Since each $AF_{ij}$ is derived from the decomposition of $MF_{i}$, there may be overlaps among the evidence $e_{ij}$. To address this issue, duplicate snippets are removed. Ultimately, $A$ is compiled as $\{E_{1},...,E_{i},...,E_{m}\}$. To intuitively demonstrate the entire procedure of our proposed method, we have summarized the process in Algorithm 1.
TABLE I: Evaluation results on API-based GPT-3.5 and GPT-4o-mini. When evaluating different LLMs, all methods also use the corresponding LLM for construction. $Attr_{r}$ and $Attr_{p}$ are used to evaluate the attribution, while Pres is used to assess the Preservation after editing. $F1_{PP}$ and $F1_{RP}$ are the harmonic means of the $Attr_{p}$ and Pres, and the $Attr_{r}$ and Pres metrics, respectively. Additionally, Average $\mathrm{F1_{PP}}$ (Ave-$\mathrm{F1_{PP}}$) and $\mathrm{F1_{RP}}$ (Ave-$\mathrm{F1_{RP}}$) are used to provide a comprehensive overview for each row. Note that the RARR paper does not provide access to the test datasets; thus, the reported results are reproduced from the publication. Bold indicates the best performance, while underlining indicates the second-best performance.
Methods GPT-3.5 GPT-4o-mini
$Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $\mathrm{F1_{PP}} Ave$-$\mathrm{F1_{RP}}$
@@
DRQA 0.424 0.647 - - - 0.265 0.576 - - - - -
EFEC 0.598 0.042 0.762 0.080 0.670 0.492 0.053 0.688 0.098 0.546 0.089 0.608
CCVER 0.624 0.066 0.928 0.123 0.747 0.563 0.051 0.931 0.097 0.702 0.110 0.728
RARR 0.649 0.058 0.868 0.109 0.743 0.493 0.080 0.904 0.147 0.638 0.128 0.691
ARE 0.670 0.756 0.910 0.826 0.772 0.623 0.722 0.902 0.802 0.737 0.814 0.755
@@
DRQA 0.431 0.673 - - - 0.251 0.544 - - - - -
EFEC 0.557 0.040 0.729 0.076 0.632 0.497 0.049 0.698 0.092 0.581 0.084 0.665
CCVER 0.630 0.069 0.937 0.129 0.753 0.634 0.045 0.933 0.086 0.755 0.108 0.754
RARR 0.646 0.060 0.850 0.112 0.734 0.508 0.058 0.899 0.109 0.649 0.111 0.692
ARE 0.716 0.807 0.914 0.857 0.803 0.634 0.755 0.881 0.813 0.737 0.835 0.770
@@
DRQA 0.237 0.490 - - - 0.080 0.222 - - - - -
EFEC 0.354 0.049 0.716 0.092 0.474 0.330 0.071 0.638 0.128 0.435 0.110 0.455
CCVER 0.372 0.047 0.932 0.089 0.532 0.341 0.056 0.961 0.106 0.503 0.098 0.518
RARR 0.356 0.097 0.862 0.174 0.504 0.292 0.079 0.916 0.145 0.443 0.160 0.474
ARE 0.463 0.559 0.899 0.689 0.611 0.413 0.469 0.904 0.618 0.567 0.654 0.589
@@
DRQA 0.127 0.283 - - - 0.122 0.308 - - - - -
EFEC 0.343 0.071 0.686 0.129 0.457 0.297 0.075 0.635 0.134 0.405 0.132 0.431
CCVER 0.292 0.071 0.967 0.132 0.449 0.310 0.062 0.980 0.117 0.471 0.221 0.460
RARR 0.340 0.078 0.904 0.144 0.494 0.269 0.064 0.944 0.120 0.419 0.132 0.457
ARE 0.412 0.438 0.917 0.593 0.569 0.387 0.414 0.914 0.570 0.544 0.582 0.557
TABLE II: Evaluation results on the open-source Llama3-70B and Llama2-70B. The metrics and other experimental settings are the same as those used in API-based evaluation. Bold indicates the best performance, while underlining indicates the second-best performance.
Methods Llama3-70B Llama2-70B
$Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $\mathrm{F1_{PP}} Ave$-$\mathrm{F1_{RP}}$
NQ
DRQA 0.382 0.620 - - - 0.483 0.522 - - - - -
EFEC 0.490 0.058 0.717 0.107 0.582 0.357 0.058 0.719 0.107 0.477 0.107 0.530
CCVER 0.597 0.071 0.921 0.132 0.724 0.423 0.068 0.813 0.126 0.539 0.129 0.632
RARR 0.646 0.060 0.850 0.112 0.734 0.516 0.066 0.594 0.119 0.552 0.115 0.643
ARE 0.682 0.739 0.898 0.811 0.759 0.584 0.682 0.926 0.785 0.716 0.798 0.738
Mintaka
DRQA 0.380 0.640 - - - 0.368 0.600 - - - - -
EFEC 0.538 0.057 0.728 0.106 0.619 0.498 0.029 0.739 0.056 0.595 0.081 0.607
CCVER 0.582 0.073 0.901 0.135 0.707 0.397 0.036 0.850 0.069 0.541 0.102 0.624
RARR 0.651 0.065 0.829 0.121 0.729 0.543 0.058 0.679 0.107 0.603 0.114 0.666
ARE 0.712 0.767 0.887 0.823 0.790 0.631 0.706 0.940 0.806 0.755 0.815 0.773
StrategyQA
DRQA 0.237 0.379 - - - 0.234 0.467 - - - - -
EFEC 0.319 0.051 0.666 0.095 0.432 0.361 0.031 0.721 0.059 0.481 0.077 0.457
CCVER 0.435 0.063 0.917 0.118 0.590 0.323 0.058 0.854 0.109 0.468 0.113 0.529
RARR 0.449 0.073 0.846 0.134 0.586 0.412 0.057 0.604 0.104 0.490 0.119 0.538
ARE 0.474 0.502 0.907 0.646 0.623 0.484 0.533 0.912 0.673 0.633 0.660 0.628
ExpertQA
DRQA 0.131 0.326 - - - 0.147 0.319 - - - - -
EFEC 0.356 0.076 0.698 0.137 0.472 0.357 0.058 0.719 0.107 0.477 0.122 0.475
CCVER 0.282 0.081 0.942 0.149 0.434 0.212 0.064 0.845 0.119 0.339 0.134 0.387
RARR 0.400 0.084 0.851 0.153 0.544 0.353 0.074 0.606 0.132 0.446 0.142 0.495
ARE 0.425 0.417 0.924 0.575 0.582 0.390 0.386 0.942 0.548 0.552 0.561 0.567
V Experiments
V-A Evaluation Setups
Dataset. We perform extensive experiments on three widely-used Question Answering (QA) datasets, as well as the AQA dataset, i.e., ExpertQA. Natural Questions (NQ) [49] is a dataset for answering questions developed by Google Search and is widely used to evaluate machine reading comprehension, information retrieval, and other tasks. StrategyQA [50] requires models to provide the reasoning process to come up with correct answers. Because, the correct answers are typically based on different paragraphs from multiple documents. Mintaka [51] is a complex multilingual dataset featuring question types such as superlative and multi-hop, which is aknowledge-intensive dataset and widely used to evaluate end-to-end question answering models. ExpertQA [52] encompasses 7 question types across 32 domains and is specifically designed to evaluate attribution-based question answering models.
@@ V-A Evaluation Setups
These datasets are used to evaluate the attribution ability in knowledge-intensive QA task. Following the standard dataset settings [9], we randomly select 150 samples from NQ, Mintaka, and StrategyQA as test datasets. For ExpertQA, we use the entire provided test set.
Baselines. The proposed ARE is compared with four post-hoc retrieval based baselines, including: DRQA [8] is a straightforward attribution model that concatenates the question and answer into a single query before retrieval.EFEC [^5] https://github.com/j6mes/acl2021-factual-error-correction [53] is an editor built on the T5 model and modifies a claim by incorporating evidence. RARR [^6] https://github.com/anthonywchen/RARR [9] first generates a series of sub-questions based on the long-form answer and retrieve evidence using these sub-questions as queries. The evidence will be used to verify the claim and further revise it by LLMs. CCVER [^7] https://github.com/jifan-chen/Fact-checking-via-Raw-Evidence [54] is a model that focusses on fact verification, it uses a few-shot approach to prompt LLMs to generate “Yes or No” type sub-questions for a given claim, which are then used as queries to retrieve evidence.
To ensure a fair comparison across different baselines, we use the same retrieval settings for EFEC as we did for RARR, and we employ RARR’s editing module for CCVER, retaining the sub-questions generation prompt of CCVER to examine the impact of generating various types of questions on attribution performance.
LLMs. We evaluated four LLMs: two open-source models (Llama2-70B-Chat, Llama3-70B-Instruct) and two API-based models (GPT-3.5-Turbo-1106, GPT-4o-mini-2024-07-18). For each specific task (e.g., long-form answer generation), the same model version was used across experiments, with variations only in prompt templates tailored to the task. The only exception is our instruction-tuned LLama3-8B-Instruct model, which is specifically fine-tuned for the fact decomposition task.
@@
Metrics. We assess the attribution and editing through several metrics: $Attr_{r}$ is a molecular-level attribution metric, which measures the recall of retrieved evidences [9]:
$Attr_{r}(X,A)=\underset{MF_{i}\in X}{\textrm{avg}}\ \underset{E_{i}\in A}{\textrm{max}}\ \mathrm{NLI}(E_{i},MF_{i}),$ (10)
@@
where, $MF_{i}\in X$ is a molecular clause and $E_{i}\in A$ is the evidence for $MF_{i}$, “max” selects the highest entailment score among all evidences, and “avg” calculates the average score of all evidence, $\mathrm{NLI}(E_{i},MF_{i})$ [^8] https://huggingface.co/google/t5_xxl_true_nli_mixture represents the model probability of $E_{i}$ entailing $MF_{i}$. The NLI model is based on T5-11B, which is trained on six datasets: MNLI, SNLI, FEVER, PAWS, SciTail, and VitaminC  [55, 56, 57, 58, 59, 60]. The input format is: “premise: {evidence} hypothesis: {sentences}”. As suggested by [9], we use the probability whose predicting result is “1” as the entailment score in the metric $Attr_{r}$.
@@
However, $Attr_{r}$ focuses solely on evidence recall, neglecting the precision of the recalled evidence. This can result in high scores when the quantity of evidence is sufficiently large. To this end, we propose $Attr_{p}$ to simultaneously assess the precision and completeness of the evidences by calculating the proportion of invalid evidence among all evidences:
$Attr_{p}=\frac{\sum_{i=1}^{m}\mathbb{I}({\textrm{NLI}_{bi}(E_{i},MF_{i})})}{m},$ (11)
where $\mathbb{I}(\mathrm{condition})$ is 1 if the condition is true, 0 otherwise; $E_{i}\in A$ and $MF_{i}\in X$. The aggregated evidence $E_{i}$ for a molecular clause $MF_{i}$ consists of multiple evidence snippets $e_{ij}$, each corresponding to an atomic fact $AF_{ij}$. We define $E_{i}$ as “invalid” if any individual snippet $e_{ij}$ fails to support its associated atomic fact.
To assess this, we use $\mathrm{NLI}_{bi}(\cdot,\cdot)$, a binary classification model based on the TRUE framework. Unlike the probabilistic version $\mathrm{NLI}(\cdot,\cdot)$, $\mathrm{NLI}_{bi}(\cdot,\cdot)$ returns “1” only if the evidence fully supports the entire sentence. This makes $Attr_{p}$ a stricter metric than $Attr_{r}$, enabling it to better evaluate the completeness of the retrieved evidence.
@@
Preservation [9] generally utilizes Levenshtein distance to measure the changed information from $X$ to $X^{\prime}$:
$Pres_{(X,X^{\prime})}=\textrm{max}(1-\frac{\textrm{Lev}(X,X^{\prime})}{\textrm{length}(X)},0),$ (12)
@@
where $Pres_{(X,X^{\prime})}$ equals 1 when $X$ is identical to $X^{\prime}$, indicating no changes. A value of 0 means $X$ and $X^{\prime}$ do not share common words, reflecting a complete divergence.
@@
To better compare with different baselines, $F1_{RP}$ [9, 10] and $F1_{PP}$ are proposed. $F1_{RP}$ and $F1_{PP}$ are calculated by the following equations:
$\mathrm{F1_{RP}=\frac{2*Attr_{r}*Pres}{Attr_{r}+Pres}},\quad\mathrm{F1_{PP}=\frac{2*Attr_{p}*Pres}{Attr_{p}+Pres}}.$ (13)